# Sharp systolic inequalities for invariant tight contact forms on principal $\mathbb{S}^{1}$-bundles over $\mathbb{S}^{2}$

### v1

@@ Abstract
The systole $\text{sys}(\alpha)$ of a contact form $\alpha$ is defined as the shortest period of closed Reeb orbits of $\alpha$. Given a non-trivial $\mathbb{S}^{1}$-principal bundle over $\mathbb{S}^{2}$ with total space $M$, we prove that the inequality $\text{sys}(\alpha)^{2}\leq\text{Vol}(M,\alpha\wedge\mathrm{d}\alpha)$ holds true for any tight contact form $\alpha$ invariant under the corresponding circle action. Moreover, equality holds exactly for Zoll contact forms, namely, contact forms whose Reeb orbits are all closed and have the same minimal period. This implies a sharp systolic inequality for rotationally symmetric Finsler metrics on $\mathbb{S}^{2}$, and a particular case of a conjecture by Viterbo.
@@ 1.2 Invariant tight contact forms and the main result
We give below examples of tight $\mathbb{S}^{1}$-invariant contact forms on bundles over $\mathbb{S}^{2}$.
@@
The main result of this work is a sharp systolic inequality for invariant tight contact forms on non-trivial $\mathbb{S}^{1}$-principal bundles over $\mathbb{S}^{2}$.
@@
If $e\neq 0$, then for all $\alpha$ in $\Omega(e)$, we have
[…]
@@ 1.3.1 Systolic inequality for Finsler metrics of revolution on $\mathbb{S}^{2}$
Remark 1.5.
The authors of [28] generalized the main result of [4] to a class of singular Riemannian metrics on $\mathbb{S}^{2}$ called rotationally symmetric spindle orbifold. In a similar fashion to Corollary 1.4, Theorem 1.2 extends the main result of [28] to rotationally symmetric Finsler metrics on spindle orbifolds, since the Reeb flow associated to such a metric lifts to a Reeb flow on $\mathbb{S}^{3}$ invariant under a free $\mathbb{S}^{1}$-action.
@@ 1.3.2 Viterbo conjecture for invariant domains
Another corollary of Theorem 1.2 is an application to a conjecture of Viterbo. Let $\omega_{0}=\sum\limits_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}$ be the standard symplectic form on $\mathbb{R}^{2n}$. We write Vol for the volume form induced by $\omega_{0}$. Viterbo stated the following conjecture in [36] (see for example [23] for an overview of the variations on this conjecture):
@@
$\text{sys}(\alpha_{U})^{2}\leq\text{Vol}(\alpha_{U})=\text{Vol}(U),$
@@ 1.4 A sketch of the proof
  • By $\mathbb{S}^{1}$-invariance of the surface of section and of the Reeb flow, studying the first-return map boils down to the study of a real-valued function, which we obtain using the decomposition of invariant contact forms obtained in Section 3.2. This function both encodes information on periodic orbits of the Reeb flow and their periods, and on the contact volume of $M$ via its integral. Building on techniques used in [4], we use this function to find a suitable $2|e|$-periodic point of the first-return map, corresponding to a periodic Reeb orbit with period less than $\sqrt{\frac{\text{Vol}(\alpha)}{|e|}}$. This yields the result.
@@ 1.5 Acknowledgments
The author is very grateful to Alberto Abbondandolo and Rémi Leclercq for numerous stimulating conversations and advice. The author is partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 281071066 – TRR 191 and by the ANR grant 21-CE40-0002 (CoSy).
@@ 2 Generalities on $\mathbb{S}^{1}$-invariant contact forms
We introduce in this section the first integral of the Reeb flow induced by the $\mathbb{S}^{1}$-symmetry, and give some of its properties. It is a general fact that to a symmetry of a dynamical system corresponds an invariant of the dynamics. In our case, this invariant is given by the moment map of the action.
@@
By $\mathbb{S}^{1}$-invariance, $\tilde{K}$ induces a function $K$ on $\mathbb{S}^{2}$. The following proposition states the standard fact that $\tilde{K}$ is indeed an invariant of the Reeb flow.
@@
which proves the first claim. We obtain from $\mathrm{d}K\circ\mathrm{d}\pi=\mathrm{d}\tilde{K}$ that $\mathrm{d}K$ vanishes at $x\in\mathbb{S}^{2}$ if and only if $\Theta_{y}$ is in the kernel of $\mathrm{d}\alpha_{y}$, with $\pi(y)=x$. Since the kernel of $\mathrm{d}\alpha$ is spanned by $R$, $\mathrm{d}K$ vanishes at $x$ if and only if $\Theta_{y}$ and $R_{y}$ are parallel. By $\mathbb{S}^{1}$-invariance of $R$ and $\Theta$, this happens if and only if $\mathbb{S}^{1}\cdot y$ and the Reeb orbit starting at $y$ coincide as set. In that case, writing $\gamma$ for the Reeb orbit starting at $y$, with minimal period $T$, and $\gamma_{\Theta}$ for $\gamma$ parameterized by the flow of $\Theta$, one has
@@ 3.1 An invariant surface of section
Let us fix $\alpha$ in $\Omega(e)$. In this section, we will construct an $\mathbb{S}^{1}$-invariant surface of section for the Reeb flow of $\alpha$. We write $Z$ for the regular level set $\tilde{K}^{-1}(0)\subset M$. Let $\mathcal{C}\subset\mathbb{S}^{2}$ be the union of the connected components of levels sets of $K$ that contain a critical point. It follows from the fact that $\mathrm{d}K$ doesn’t vanish on $\mathbb{S}^{2}\setminus\mathcal{C}$, that the difference $\mathbb{S}^{2}\setminus\mathcal{C}$ is a union of cylinders $(\Gamma_{i})_{i\in I}$. By Remark 2.5, $\pi(Z)$ is connected. Hence, there is a unique index $i_{0}$ such that $K$ has zeroes on $\Gamma_{i_{0}}$.
@@
Writing $\tilde{\gamma}_{i}$ for a lift of $\gamma_{i}$ to $M$, we set $\Sigma_{i}$ as the image of $\tilde{\gamma}_{i}$ under the $\mathbb{S}^{1}$ action, i.e. $\Sigma_{i}=\mathbb{S}^{1}\cdot\tilde{\gamma}_{i}=\pi^{-1}(\gamma_{i})$. Their union $\Sigma=\bigcup\limits_{i\in I}\Sigma_{i}$ is a disjoint union of cylinders. Its boundary components have the form $\mathbb{S}^{1}\cdot x$, with $x$ a critical point of $K$, and as such are both a closed Reeb orbit and an $\mathbb{S}^{1}$-orbit. The following proposition states that $\Sigma$ is a surface of section.
@@ 3.2 A decomposition for invariant contact forms
Since $\mathbb{S}^{2}\setminus\pi(Z)$ is the union of two disks, and because the only $\mathbb{S}^{1}$-principal bundle over a disk is the trivial one, each of the two components of $M\setminus Z$ is a solid torus, the fundamental group of which being generated by the $\mathbb{S}^{1}$-orbits of the action. Let $\sigma$ be an $\mathbb{S}^{1}$-invariant closed one-form on $M\setminus Z$ such that $\int_{\mathbb{S}^{1}\cdot x}\sigma=1$ for all $x$ in $M\setminus Z$, the $\mathbb{S}^{1}\cdot x$ being oriented by the action. We will occasionally write $\sigma^{+}$ (resp. $\sigma^{-}$) for the restriction of $\sigma$ to $(K>0)$ (resp. to ($K<0$)).
@@
where $\beta$ is a one-form on $\mathbb{S}^{2}\setminus\pi(Z)$.
@@
We write $D_{1}$ (resp. $D_{2}$) for the 2-disk $(K>0)$ (resp. $(K<0)$), and $\text{Crit}(f)$ for the set of critical values of a real-valued function $f$. The bundle $M$ is obtained by gluing the two bundles $T_{1}=\pi^{-1}(D_{1})$ and $T_{2}=\pi^{-1}(D_{2})$ via the attaching map described above. To understand the attaching map, we must understand the orientations of $\partial T_{1}$ and $\partial T_{2}$. Let $Y$ be an $\mathbb{S}^{1}$-invariant gradient-like vector field for $\tilde{K}$ on $M$, i.e. a vector field vanishing exactly at critical points of $\tilde{K}$, and such that $\mathrm{d}\tilde{K}(Y)>0$ otherwise. Since $\alpha\wedge\mathrm{d}\alpha(R,Y,X)=\mathrm{d}\alpha(Y,X)=\mathrm{d}K(Y)>0$ on $M\setminus\text{Crit}(\tilde{K})$, $(R,Y,X)$ is a positive basis on $M\setminus\text{Crit}(\tilde{K})$. Thus, $(\pi_{*}R,\pi_{*}Y)$ is a positive basis on $\mathbb{S}^{2}\setminus\text{Crit}(K)$. We get that the induced boundary orientation of $\partial D_{1}$ (resp. of $\partial D_{2}$) coincides with (resp. is the opposite of) the orientation induced by the Reeb flow. Following Remark 3.3, the attaching map $\phi$ acts on the level of fundamental groups via the map $\phi_{*}=\begin{pmatrix}1&-e\\ 0&-1\end{pmatrix}$ in the basis $(h_{1},\mu_{1})$ and $(h_{2},\mu_{2})$. Hence, the induced map $\phi^{*}\colon H^{1}(\partial T_{2},\mathbb{Z})\rightarrow H^{1}(\partial T_{1},\mathbb{Z})$ is $\begin{pmatrix}1&0\\ -e&-1\end{pmatrix}$ in the dual basis $(h_{1}^{*},\mu_{1}^{*})$ and $(h_{2}^{*},\mu_{2}^{*})$. Since $\sigma^{+}=h_{1}^{*}$ on $T_{1}$ and $\sigma^{-}=h_{2}^{*}$ on $T_{2}$, and since $\pi(c_{i_{0},0})=\mu_{1}$ in $\pi_{1}(\partial T_{1})$, we get
@@ 3.3 A family of potentials
Since the curves $\pi(c_{i,k})$ are closed curves that uniformly converge to $\epsilon_{i}^{\pm}$ when $k$ approaches to $K_{i}^{\pm}$, $J_{i}$ can be extended continuously to $[K^{-}_{i},K^{+}_{i}]$. Moreover, since $\lim\limits_{k\rightarrow 0,k>0}J_{i_{0}}(k)=\lim\limits_{k\rightarrow 0,k>0}\int_{c_{i,k}}\alpha-k\sigma=\int_{c_{i_{0},0}}\alpha$, and similarly $\lim\limits_{k\rightarrow 0,k<0}J_{i_{0}}(k)=\int_{c_{i_{0},0}}\alpha$, $J_{i_{0}}$ is continuous at 00.
@@
An immediate corollary of Lemma 3.5 is the below identity relating $\tau$ and $J_{i}$.
@@
We now give some properties of the family of potentials $(J_{i})_{i\in I}$. First of all, Lemma 3.8 tells us that the integrals of the $(J_{i})_{i\in I}$ are related to the volume.
@@
This follows immediately from Corollary 3.6 in addition to the positivity of $\tau$. ∎
@@
Without loss of generality, we can assume that $f(b)\geq 0$ and $0<a<b$. If there is an $x_{0}\in]a,b[$ such that $f(x_{0})\leq 0$, it follows immediately from the inequality (1), plus the fact that $x_{0}>0$, that $f^{\prime}(x_{0})<0$. Furthermore, the inequality (1) implies that $f(x)>0$ whenever $f^{\prime}(x)=0$. Thus we get that $f^{\prime}$ is negative on $[x_{0},b[$. As a consequence, $f(x)<0$ on $[x_{0},b]$ and in particular $f(b)<0$, which is a contradiction. ∎
@@
[…]
@@
Let $T$ be a component of $M\setminus Z$. Assume without loss of generality that $\tilde{K}$ is positive on $T$. We start by proving the result when $K$ has isolated critical points.
@@
    • $\tilde{g}$ is constant equal to one on $\tilde{K}^{-1}([k-\eta,k+\eta])$;
@@
4 Proof of Theorem 1.2
In this section, we prove Theorem 1.2. Since the case of $\mathbb{S}^{1}$-actions transverse to the contact structure has been taken care of in Proposition 2.4, we focus now on the case where $e>0$ and for which the set of set of Legendrian $\mathbb{S}^{1}$-orbits is non-empty and connected .

### v2

@@ Abstract
The systole of a contact form $\alpha$ is defined as the shortest period of closed Reeb orbits of $\alpha$. Given a non-trivial $\mathbb{S}^{1}$-principal bundle over $\mathbb{S}^{2}$ with total space $M$, we prove a sharp systolic inequality for the class of tight contact form on $M$ invariant under the $\mathbb{S}^{1}$-action. This inequality exhibits a behavior which depends on the Euler class of the bundle in a subtle way. As applications, we prove a sharp systolic inequality for rotationally symmetric Finsler metrics on $\mathbb{S}^{2}$, a systolic inequality for the shortest contractible closed Reeb orbit, and a particular case of a conjecture by Viterbo.
@@ 1.2 Invariant tight contact forms and the main result
The classification of invariant contact structures on principal $\mathbb{S}^{1}$-bundles have been achieved in [35] and [27], see also [25, Section 5] for a detailed discussion of $\mathbb{S}^{1}$-invariant contact structures on $\mathbb{S}^{3}$. We give below examples of tight $\mathbb{S}^{1}$-invariant contact forms on bundles over $\mathbb{S}^{2}$.
@@
The main result of this work is a sharp systolic inequality for invariant tight contact forms on non-trivial $\mathbb{S}^{1}$-principal bundles over $\mathbb{S}^{2}$. This inequality exhibits a behavior which depends on the Euler number of the bundle in a non-trivial way.
@@
  • If $e<0$ or if $e\in\{1,2\}$, then for all $\alpha$ in $\Omega(e)$, we have
    $\text{sys}(\alpha)^{2}\leq\frac{1}{|e|}\text{Vol}(\alpha).$
@@
  • If $e>2$, then for all $\alpha$ in $\Omega(e)$, we have
    $\text{sys}(\alpha)^{2}<\frac{1}{2}\text{Vol}(\alpha).$
    Moreover, the supremum of $\frac{\text{sys}(\alpha)^{2}}{\text{Vol}(\alpha)}$ on $\Omega(e)$ is $\frac{1}{2}$ (and is not achieved).
@@
  • The fact that the supremum of $\frac{\text{sys}(\alpha)^{2}}{\text{Vol}(\alpha)}$ on $\Omega(e)$ is not achieved when $e>2$ can be understood as follows [^1] The author wishes to thank Christian Lange for pointing out that the systolic inequality had to be different for Euler numbers greater than 2.: the classification of Zoll contact forms in dimension 3, achieved in [14], implies that there is no positive Zoll contact form on $M$ when $e>2$. Moreover, contact forms that are local maximizers of the systolic ratio $\frac{\text{sys}(\alpha)^{2}}{\text{Vol}(\alpha)}$ in the class of $\mathbb{S}^{1}$-invariant contact forms are necessarily Zoll, as proved in [38, Theorem 3.4]. Thus the equality cannot be reached for the systolic inequality for invariant forms on $M$ when $e>2$.
@@ 1.3.1 Systolic inequality for Finsler metrics of revolution on $\mathbb{S}^{2}$
1.3.2 Contractible systole and Besse contact forms
Another class of special contact forms, generalizing the Zoll ones, is the class of Besse contact forms. A contact form is Besse when its Reeb flow only has periodic orbits (not necessarily of same minimal period as in the Zoll case). A classical result due to Wadsley [44], or in the three dimensional case an earlier result of Epstein [22], implies that the Reeb flow of a Besse contact form is periodic. The theorem of Epstein implies that all orbits but finitely many ones of the Reeb flow of a Besse form have the same minimal period $T$, and the minimal period of the finitely many others divides $T$. Similarly to the principle conjectured in [38] and proved in [6], [14], [2], that Zoll contact forms are exactly the local maximizers of the systolic ratio, some variations of the systolic ratio were studied in [8], [12] and [34] for which Besse contact forms are exactly the local maximizers. In [34], the authors study the contractible systole of a contact form $\alpha$, denoted $\text{sys}_{\text{contr}}(\alpha)$, namely the shortest period of a contractible periodic orbit of the Reeb flow of $\alpha$. They prove a systolic inequality for the contractible systole for the contact forms induced by a class of singular Riemannian metrics on $\mathbb{S}^{2}$ called rotationally symmetric spindle orbifolds (Theorem A in [34]), generalizing the main result of [4]. Moreover, equality holds exactly when the spindle orbifold is Besse.
Theorem 1.2 also implies a systolic inequality for the contractible systole, stated in Corollary 1.5 below, whose proof will be given in 4.2. It recovers as a particular case Theorem A in [34]
Corollary 1.5.
Let $e\neq 0$, then for all $\alpha\in\Omega(e)$,
$\text{sys}_{\text{contr}}(\alpha)^{2}\leq|e|\text{Vol}(\alpha).$
Moreover, equality holds if and only if $\alpha$ is Zoll when $e\leq 2$, and if and only if $\alpha$ is Besse with two singular orbits of same minimal period, when $e>2$. In the latter case, the singular orbits have period $\frac{T}{e}$ if $e$ is odd, $\frac{2T}{e}$ is $e$ is even, where $T$ is the minimal period of the regular orbits.
@@ 1.3.3 Viterbo conjecture for invariant domains
Another corollary of Theorem 1.2 is an application to a conjecture of Viterbo. Let $\omega_{0}=\sum\limits_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}$ be the standard symplectic form on $\mathbb{R}^{2n}$. We write $\text{Vol}(U,\omega_{0}^{n})$ for the symplectic volume of a set $U$. Viterbo stated the following conjecture in [43] (see for example [29] for an overview of the variations on this conjecture):
@@
$\text{sys}(\alpha_{U})^{2}\leq\text{Vol}(\alpha_{U})=\text{Vol}(U,\omega_{0}^{n}),$
@@ 1.4 A sketch of the proof
  • By $\mathbb{S}^{1}$-invariance of the surface of section and of the Reeb flow, studying the first-return map boils down to the study of a real-valued function, which we obtain using the decomposition of invariant contact forms obtained in Section 3.2. This function both encodes information on periodic orbits of the Reeb flow and their periods, and on the contact volume of $M$ via its integral. Building on techniques used in [4], we use this function to find a suitable periodic point of the first-return map, corresponding to a periodic Reeb orbit with controlled period. This yields the result.
@@ 1.5 Acknowledgments
The author is very grateful to Alberto Abbondandolo and Rémi Leclercq for numerous stimulating conversations and advice, as well as to Christian Lange for his careful reading of the first version of this work. The author is partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 281071066 – TRR 191 and by the ANR grant 21-CE40-0002 (CoSy).
@@ 2 Generalities on $\mathbb{S}^{1}$-invariant contact forms
We introduce in this section the first integral of the Reeb flow induced by the $\mathbb{S}^{1}$-symmetry, and give some of its properties. It is a general fact that to a symmetry of a dynamical system corresponds an invariant of the dynamics. In our case, this invariant is given by the moment map of the action. Recall that $\Omega(e)$ is the set on $\mathbb{S}^{1}$-invariant positive tight contact forms on the total space of the principal $\mathbb{S}^{1}$-bundle over $\mathbb{S}^{2}$ with Euler number $e$.
@@
By $\mathbb{S}^{1}$-invariance, $\tilde{K}$ induces a function on $\mathbb{S}^{2}$, which we denote $K$.
@@
which proves the first claim. We obtain from $\mathrm{d}K\circ\mathrm{d}\pi=\mathrm{d}\tilde{K}$ that $\mathrm{d}K$ vanishes at $x\in\mathbb{S}^{2}$ if and only if $\Theta_{y}$ is in the kernel of $\mathrm{d}\alpha_{y}$, with $\pi(y)=x$. Since the kernel of $\mathrm{d}\alpha$ is spanned by $R$, $\mathrm{d}K$ vanishes at $x$ if and only if $\Theta_{y}$ and $R_{y}$ are parallel. By $\mathbb{S}^{1}$-invariance of $R$ and $\Theta$, this happens if and only if $\mathbb{S}^{1}\cdot y$ and the Reeb orbit starting at $y$ coincide as sets. In that case, writing $\gamma$ for the Reeb orbit starting at $y$, with minimal period $T$, and $\gamma_{\Theta}$ for $\gamma$ parameterized by the flow of $\Theta$, one has
@@ 3.1 An invariant surface of section
Let us fix $\alpha$ in $\Omega(e)$. In this section, we will construct an $\mathbb{S}^{1}$-invariant surface of section for the Reeb flow of $\alpha$. We write $Z$ for the regular level set $\tilde{K}^{-1}(0)\subset M$. Let $\mathcal{C}\subset\mathbb{S}^{2}$ be the union of the connected components of levels sets of $K$ that contain a critical point. It follows from the fact that $\mathrm{d}K$ does not vanish on $\mathbb{S}^{2}\setminus\mathcal{C}$, that $\mathbb{S}^{2}\setminus\mathcal{C}$ decomposes as a union of cylinders $(\Gamma_{i})_{i\in I}$. By Remark 2.5, $\pi(Z)$ is connected. Hence, there is a unique index $i_{0}$ such that $K$ has zeroes on $\Gamma_{i_{0}}$.
@@
Writing $\tilde{\gamma}_{i}$ for a lift of $\gamma_{i}$ to $M$, we set $\Sigma_{i}$ as the image of $\tilde{\gamma}_{i}$ under the $\mathbb{S}^{1}$-action, i.e. $\Sigma_{i}=\mathbb{S}^{1}\cdot\tilde{\gamma}_{i}=\pi^{-1}(\gamma_{i})$. The $\Sigma_{i}$ are disjoint cylinders, whose boundary components have the form $\mathbb{S}^{1}\cdot x$, with $x$ a critical point of $K$, and as such are both closed Reeb orbits and $\mathbb{S}^{1}$-orbits. We write $\Sigma$ for their union $\bigcup\limits_{i\in I}\Sigma_{i}$. The following proposition states that $\Sigma$ is a surface of section.
@@ 3.2 A decomposition for invariant contact forms
Since $\mathbb{S}^{2}\setminus\pi(Z)$ is the union of two disks, and because the only $\mathbb{S}^{1}$-principal bundle over a disk is the trivial one, each of the two components of $M\setminus Z$ is a solid torus, the fundamental group of which being generated by the $\mathbb{S}^{1}$-orbits of the action. Let $\sigma$ be an $\mathbb{S}^{1}$-invariant closed one-form on $M\setminus Z$ such that $\int_{\mathbb{S}^{1}\cdot x}\sigma=1$ for all $x$ in $M\setminus Z$, the orientation of $\mathbb{S}^{1}\cdot x$ being given by the action. We will occasionally write $\sigma^{+}$ (resp. $\sigma^{-}$) for the restriction of $\sigma$ to $(K>0)$ (resp. to ($K<0$)).
@@
for some one-form $\beta$ on $\mathbb{S}^{2}\setminus\pi(Z)$.
@@
We write $D_{1}$ (resp. $D_{2}$) for the 2-disk $(K>0)$ (resp. $(K<0)$), and $\text{Crit}(f)$ for the set of critical values of a real-valued function $f$. The bundle $M$ is obtained by gluing the two bundles $T_{1}=\pi^{-1}(D_{1})$ and $T_{2}=\pi^{-1}(D_{2})$ via the attaching map described above. To understand the attaching map, we must understand the orientations of $\partial T_{1}$ and $\partial T_{2}$. Let $Y$ be an $\mathbb{S}^{1}$-invariant gradient-like vector field for $\tilde{K}$ on $M$, i.e. a vector field vanishing exactly at critical points of $\tilde{K}$, and such that $\mathrm{d}\tilde{K}(Y)>0$ otherwise. Since $\alpha\wedge\mathrm{d}\alpha(R,Y,\Theta)=\mathrm{d}\alpha(Y,\Theta)=\mathrm{d}K(Y)>0$ on $M\setminus\text{Crit}(\tilde{K})$, $(R,Y,\Theta)$ is a positive basis on $M\setminus\text{Crit}(\tilde{K})$. Thus, $(\pi_{*}R,\pi_{*}Y)$ is a positive basis on $\mathbb{S}^{2}\setminus\text{Crit}(K)$. We get that the induced boundary orientation of $\partial D_{1}$ (resp. of $\partial D_{2}$) coincides with (resp. is the opposite of) the orientation induced by the Reeb flow. Following Remark 3.3, the attaching map $\phi$ acts on the level of fundamental groups via the map $\phi_{*}=\begin{pmatrix}1&-e\\ 0&-1\end{pmatrix}$ in the basis $(h_{1},\mu_{1})$ and $(h_{2},\mu_{2})$. Hence, the induced map $\phi^{*}\colon H^{1}(\partial T_{2},\mathbb{Z})\rightarrow H^{1}(\partial T_{1},\mathbb{Z})$ is $\begin{pmatrix}1&0\\ -e&-1\end{pmatrix}$ in the dual basis $(h_{1}^{*},\mu_{1}^{*})$ and $(h_{2}^{*},\mu_{2}^{*})$. Since $\sigma^{+}=h_{1}^{*}$ on $T_{1}$ and $\sigma^{-}=h_{2}^{*}$ on $T_{2}$, and since $\pi(c_{i_{0},0})=\mu_{1}$ in $\pi_{1}(\partial T_{1})$, we get
@@ 3.3 A family of potentials
Since the curves $\pi(c_{i,k})$ are closed curves that uniformly converge to $\epsilon_{i}^{\pm}$ when $k$ goes to $K_{i}^{\pm}$, $J_{i}$ can be extended continuously to $[K^{-}_{i},K^{+}_{i}]$. Moreover, since $\lim\limits_{k\rightarrow 0,k>0}J_{i_{0}}(k)=\lim\limits_{k\rightarrow 0,k>0}\int_{c_{i,k}}\alpha-k\sigma=\int_{c_{i_{0},0}}\alpha$, and similarly $\lim\limits_{k\rightarrow 0,k<0}J_{i_{0}}(k)=\int_{c_{i_{0},0}}\alpha$, $J_{i_{0}}$ is continuous at 00.
@@
An immediate corollary of Lemma 3.5 is the identity below relating $\tau$ and $J_{i}$.
@@
We now give some properties of the family of potentials $(J_{i})_{i\in I}$. First of all, Lemma 3.8 tells us how the integral of the $J_{i}$’s are related to the volume.
@@
This follows immediately from Lemma 3.5 and Corollary 3.6 in addition to the positivity of $\tau$. ∎
@@
Without loss of generality, we can assume that $f(b)\geq 0$ and $0<a<b$. If there exists $x_{0}\in]a,b[$ such that $f(x_{0})\leq 0$, it follows immediately from inequality (1) at $x_{0}$, plus the fact that $x_{0}>0$, that $f^{\prime}(x_{0})<0$. Furthermore, the inequality (1) implies that $f(x)>0$ whenever $f^{\prime}(x)=0$. Thus $f^{\prime}$ is negative on $[x_{0},b[$. As a consequence, $f(x)<0$ on $[x_{0},b]$ and in particular $f(b)<0$, which is a contradiction. ∎
@@
$\begin{array}[]{ccccc}f&\colon&[\min(K),\max(K)]\setminus\text{Crit}(K)&\rightarrow&\mathbb{R}\\ &&k&\mapsto&\int_{K^{-1}(k)}\beta\end{array}$
@@
Let $T$ be a component of $M\setminus Z$. Assume without loss of generality that $\tilde{K}$ is positive on $T$.
@@
    • $\tilde{g}$ is constant equal to $1$ on $\tilde{K}^{-1}([k-\eta,k+\eta])$;
@@
4 Proof of the main results
4.1 Proof of Theorem 1.2
In this section, we prove Theorem 1.2. Since the case of $\mathbb{S}^{1}$-actions transverse to the contact structure has been taken care of in Proposition 2.4, we focus now on the case where $e>0$ and for which the set of Legendrian $\mathbb{S}^{1}$-orbits is non-empty and connected. We prove separately in Proposition 4.1 the optimality of the constant $\frac{1}{2}$ in the systolic inequality when the Euler number is larger than $2$.
Proposition 4.1.
If $e>2$, then the supremum of $\frac{\text{sys}(\alpha)^{2}}{\text{Vol}(\alpha)}$ on $\Omega(e)$ is at least $\frac{1}{2}$.